5√(2)+6√(8)-7√(45)
Answer by RAY100(1637)

5sqrt2 +6sqrt8 - 7sqt45
.
5 sqrt2 +6 sqrt(4*2) -7sqrt(9*5)
.
5sqrt2 +6{sqrt4*sqrt2} -7{sqrt9*sqrt5}
.
5sqrt2 +6*2sqrt2 -7*3sqrt5
.
5sqrt2 +12sqrt2 -21sqrt5
.
17sqrt2 -21sqrt5
.
check
5sqrt2 +6sqrt8 -7sqrt45 = 17sqrt2 -21sqrt5
.
7.07 + 16.97 -46.96 = 24.04 -46.96
.
-22.92 = -22.92,,,,,,,ok